The circumference of a circle is given by the formula C = 2πr. Find r if the circumference C is 176 meters and π is 22/7.
? meters

Answer by ewatrrr(24785)


Hi
C = 2πr OR r = (1/2)7/22C
C = 176m
r = (1/2)7/22*176m = 28m
CHECKING our Answer***
2*22/7*28m = 176m